# Where is the Super-Virial Gas? The Supply from hot inflows

### v1

@@ 5 Conclusions
Several observations have recently established that the MW contains gas at temperatures higher than its virial temperature. However, location, origin and properties of this gas are still unknown. In this paper, we analyze a suite of idealized simulations to investigate this newly-discovered gas phase and find this super-virial phase gas in our simulations of a MW-like galaxy. Our main findings are the following:
Most of the super-virial gas is within 20 kpc from the center of the galaxy (Left Panel of Figure 1), which implies they are situated mostly in extended disc region, not in the CGM.
The mass of this gas is roughly constant, in the range . This phase is retained in the simulations even at 1.5 Gyr (Figure 2).
We also investigate the effect of stellar feedback in heating up the gas. We see no change in the gas mass in the case of no feedback run (Figure 2). However, we see that some cold K gas in the central galaxy is heated up to super-virial phase which is absent in no feedback run. However, the feedback heating contributes only less than 1% of the super-virial gas, implying this mechanism of heating is sub-dominant, at least for the FIRE-feedback prescription (Figure 5).
We also explore effect of CR in the heating of the gas. With high and low CR initial setup, we found no change in the super virial gas mass hence implying that the CR has little role in heating up the gas to super-virial temperatures (Figure 2).
While investigating the origin of super-virial gas, we found that the dominant contribution to heating is caused by a rotating inflow of the virial CGM phase as in the model of Stern et al. (2024). In this model gravitational energy is converted to heat mainly via compressive heating in the inflow. We show this process causes the 1-2% of gas infalling closest to the rotation axis to reach super-virial temperatures just before cooling and joining the disk (Figs. 7, 8). Gas inflowing near the rotation axis has a lower density and thus suffers less cooling losses and is more compressively heated than gas farther away from the axis, allowing it to reach higher temperatures. Gas stripped from satellites also inflows onto central galaxy and heats up similarly (Figure 6).
In conclusion, we identify compressive heating of the hot inflow as a significant contributor to the super-virial gas in the Milky Way. However, several caveats require further investigation in future studies, including more detailed comparisons with observational data (see Section 4).
@@ Acknowledgments
We also thank FIRE collaboration for useful discussions and suggestions. MR acknowledges support from the Center for Cosmology and particle physics fellowship at Ohio State University. MR acknowledges ACCESS allocations PHY240003. MR also acknowledges Aspen Center for Physics and Simons Foundation as this work was performed in part at the Aspen Center for Physics, which is supported by a grant from the Simons Foundation (1161654, Troyer). KS acknowledges support from the Black Hole Initiative at Harvard University, which is funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation, and acknowledges ACCESS allocations TG-PHY220027 and TG-PHY220047 and Frontera allocation AST22010. JS was supported by the Israel Science Foundation (grant No. 2584/21). The computations in this work were run at facilities supported by the Bridges, University of Pittsburgh.

### v2

@@ 5 Conclusions
Several observations have recently established that the MW contains gas at temperatures higher than its virial temperature. However, the location, origin, and properties of this gas are still unknown. In this paper, we analyze a suite of idealized simulations to investigate this newly-discovered gas phase and find this super-virial phase gas in our simulations of a MW-mass galaxy. Our main findings are the following:
The mass of this gas phase is roughly constant over time, in the range , but does not significantly contribute to the CGM mass budget. This phase is retained in the simulations even at 1.5 Gyr (Figure 2).
We also investigate the effect of stellar feedback in heating the gas. The feedback heating contributes only less than 1% of the super-virial gas, implying the mechanism of feedback heating is sub-dominant, at least for the FIRE-feedback prescription (Figure 5).
We also explore the effect of CR in the heating of the gas. With high and low CR initial setups, we found minimal differences in the super virial gas mass budget, however, the gas mass in CR runs shows more fluctuation, caused by additional CR feedback effect. Therefore, it implies that the CR has little role in heating the gas to super-virial temperatures (Figure 2).
While investigating the origin of super-virial gas, we found that the dominant contribution to heating is caused by the adiabatic compression and shocks of the rotating inflow of the virial CGM phase, as in the model of Stern et al. (2024). In this model, gravitational energy is converted to heat mainly via compressive heating in the inflow. Later on, they are shock-heated before settling onto the disc, indicated by the entropy jump (bottom panel of Figure 3). We show this process causes the % of infalling virial gas, closest to the rotation axis, to reach super-virial temperatures just before cooling and joining the disk (Figures 7). Gas inflowing near the rotation axis has a lower density and thus suffers less cooling losses and is more compressively heated than gas farther away from the axis, allowing it to reach higher temperatures. Gas stripped from satellites also inflows onto the central galaxy and heats up similarly (Figure 6).
In conclusion, we identify a combination of compressive heating and shock-heating of the hot inflow as a significant contributor to the super-virial gas phase in our MW-mass galaxy simulation using FIRE-2 physics. However, this process is not the only channel, but rather one of the several channels by which super-virial gas can be formed. Other channels include feedback heating and CR heating, which are not dominant in our simulation and are explored in other studies. In reality, the combination of all of these processes can give rise to the super-virial phase in a galaxy.
@@ Acknowledgments
We thank Todd Thompson and the FIRE collaboration for useful discussions and suggestions. MR acknowledges support from the CCAPP fellowship at The Ohio State University and ACCESS allocations of PHY240003. MR acknowledges the Aspen Center for Physics and Simons Foundation, as part of this work was performed there, which is supported by a grant from the Simons Foundation (1161654, Troyer). KS acknowledges support from the Black Hole Initiative at Harvard University, which is funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation and acknowledges ACCESS allocations TG-PHY220027 and TG-PHY220047 and Frontera allocation AST22010. SM is grateful for the grant provided by the National Aeronautics and Space Administration (NASA) through Chandra Award Number AR0-23014X issued by the Chandra X-ray Center, which is operated by the Smithsonian Astrophysical Observatory for and on behalf of the National Aeronautics Space Administration under contract NAS8-03060. S.M. is grateful for the NASA ADAP grant 80NSSC22K1121. JS was supported by the Israel Science Foundation (grant No. 2584/21). The computations in this work were run at facilities supported by the Bridges, University of Pittsburgh.
Appendix A Additional Figures
We show the subset of individual particle tracking for set-1 particles in Figure A.1, for which we show the median with percentile error in the main text. We note that most of the particles behave similarly like the median trajectory within the error bar. Therefore, we think that the median with error is pretty representative of the whole sample. It is also true for the set-2 and set-3 particles, which we don’t show here.